# LoCoDL: Communication-Efficient Distributed Learning with Local Training and Compression

### v1

@@ 4 Experiments
We evaluate the performance of our proposed method LoCoDL and compare it with several other methods that also allow for CC and converge linearly to $x^{\star}$. We also include GradSkip (Maranjyan et al., 2023) and Scaffold (McMahan et al., 2017) in our comparisons for benchmarking purposes. We focus on a regularized logistic regression problem, which has the form (1) with
@@
We considered several datasets and show here the results with the ‘diabetes’ and ‘a5a’ datasets from the LibSVM library (Chang & Lin, 2011). Additional experiments are shown in the Appendix. We prepared each dataset by first shuffling it, then distributing it equally among the $n$ clients (since $m$ in (14) is an integer, the remaining datapoints were discarded).
We used four different compression operators in the class $\mathbb{U}(\omega)$, for some $\omega\geq 0$:
  1. 1.
    rand-$k$ for some $k\in[d]$, which communicates $32k+k\lceil\log_{2}(d)\rceil$ bits. Indeed, the $k$ randomly chosen values are sent in the standard 32-bits IEEE floating-point format, and their locations are encoded with $k\lceil\log_{2}(d)\rceil$ additional bits. We have $\omega=\frac{d}{k}-1$.
  2. 2.
    Natural Compression (Horváth et al., 2022), a form of quantization in which floats are encoded into 9 bits instead of 32 bits. We have $\omega=\frac{1}{8}$.
  3. 3.
    A combination of rand-$k$ and Natural Compression, in which the $k$ chosen values are encoded into 9 bits, which yields a total of $9k+k\lceil\log_{2}(d)\rceil$ bits. We have $\omega=\frac{9d}{8k}-1$.
  4. 4.
    The $l_{1}$-selection compressor, defined as $C(x)=\text{sign}(x_{j})\|x\|_{1}e_{j}$, where $j$ is chosen randomly in $[d]$, with the probability of choosing $j^{\prime}\in[d]$ equal to $|x_{j^{\prime}}|/\|x\|_{1}$, and $e_{j}$ is the $j$-th standard unit basis vector in $\mathbb{R}^{d}$. $\text{sign}(x_{j})\|x\|_{1}$ is sent as a 32-bits float and the location of $j$ is indicated with $\lceil\log_{2}(d)\rceil$, so that this compressor communicates $32+\lceil\log_{2}(d)\rceil$ bits. Like with rand-$1$, we have $\omega=d-1$.
@@
The results are shown in Figures 1 and 2. We can see that LoCoDL, when combined with rand-$k$ and Natural Compression, converges faster than all other algorithms, with respect to the total number of communicated bits per client. We chose two different numbers $n$ of clients, one with $n<d$ and another one with $n>2d$, since the compressor of CompressedScaffnew is different in the two cases $n<2d$ and $n>2d$ (Condat et al., 2022a). LoCoDL outperforms CompressedScaffnew in both cases. As expected, all methods exhibit faster convergence with larger $n$. Remarkably, ADIANA, which has the best theoretical complexity for large $n$, improves upon DIANA but is not competitive with the LT-based methods CompressedScaffnew, 5GCS-CC, and LoCoDL. This illustrates the power of doubly-accelerated methods based on a successful combination of LT and CC. In this class, our new proposed LoCoDL algorithm shines. For all algorithms, we used the theoretical parameter values given in their available convergence results (Corollary 3.2 for LoCoDL). We tried to tune the parameter values, such as $k$ in rand-$k$ and the (average) number of local steps per round, but this only gave minor improvements. For instance, ADIANA in Figure 2 was a bit faster with the best value of $k=20$ than with $k=30$. Increasing the learning rate $\gamma$ led to inconsistent results, with sometimes divergence.
The results of additional experiments on the ‘w1a’ and ‘australian’ datasets from the LibSVM library (Chang & Lin, 2011), in the same setting, are shown in Figures 3 and 4. Consistent with our previous findings, LoCoDL outperforms the other algorithms in terms of communication efficiency.
@@
5 Conclusion
We have proposed LoCoDL, which combines a probabilistic Local Training mechanism similar to the one of Scaffnew and Communication Compression with a large class of unbiased compressors. This successful combination makes LoCoDL highly communication-efficient, with a doubly accelerated complexity with respect to the model dimension $d$ and the condition number of the functions. In practice, LoCoDL outperforms other algorithms, including ADIANA, which has an even better complexity in theory obtained from Nesterov acceleration and not Local Training. This again shows the relevance of the popular mechanism of Local Training, which has been widely adopted in Federated Learning. A venue for future work is to implement bidirectional compression (Liu et al., 2020; Philippenko & Dieuleveut, 2021). We will also investigate extensions of our method with calls to stochastic gradient estimates, with or without variance reduction, as well as partial participation. These two features have been proposed for Scaffnew in Malinovsky et al. (2022) and Condat et al. (2023), but they are challenging to combine with generic compression.
Acknowledgement
This work was supported by the SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence (SDAIA-KAUST AI).
@@ Acknowledgement
$=\left\|S\mathbf{\hat{x}}^{t}-\mathbf{x}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}+2n\mathbb{E}\!\left[\left\|\bar{d}^{t}-\mathbb{E}\!\left[\bar{d}^{t}\;|\;\mathcal{F}^{t},\theta=1\right]\right\|^{2}\;|\;\mathcal{F}^{t},\theta=1\right]$
@@
$+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left\|\mathbf{u}^{t}-\mathbf{u}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}+\frac{p\chi}{\gamma}\left\|W\mathbf{\hat{x}}^{t}\right\|^{2}_{\boldsymbol{\mathcal{X}}}-2\langle\mathbf{u}^{t}-\mathbf{u}^{\star},\mathbf{\hat{x}}^{t}\rangle_{\boldsymbol{\mathcal{X}}}$
$=\frac{1}{\gamma}\left\|\mathbf{w}^{t}-\mathbf{w}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left(1-\frac{p^{2}\chi}{1+2\omega}\right)\left\|\mathbf{u}^{t}-\mathbf{u}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}-\frac{p}{\gamma}\big{(}2\rho-\rho^{2}(1+\omega_{\mathrm{av}})-\chi\big{)}\left\|W\mathbf{\hat{x}}^{t}\right\|^{2}_{\boldsymbol{\mathcal{X}}}.$

### v2

@@ 4 Experiments
We evaluate the performance of our proposed method LoCoDL and compare it with several other methods that also allow for CC and converge linearly to $x^{\star}$. We also include GradSkip (Maranjyan et al., 2022) and Scaffold (McMahan et al., 2017) in our comparisons. We focus on a regularized logistic regression problem, which has the form (1) with
@@
We considered several datasets from the LibSVM library (Chang & Lin, 2011) (3-clause BSD license). We show the results with the ‘a5a’ , ‘diabetes’, ‘w1a’ datasets in Figures 1, 2, 3, respectively. Other datasets are shown in the Appendix. We prepared each dataset by first shuffling it, then distributing it equally among the $n$ clients (since $m$ in (11) is an integer, the remaining datapoints were discarded). We used four different compression operators in the class $\mathbb{U}(\omega)$, for some $\omega\geq 0$:
$\bullet$  rand-$k$ for some $k\in[d]$, which communicates $32k+k\lceil\log_{2}(d)\rceil$ bits. Indeed, the $k$ randomly chosen values are sent in the standard 32-bits IEEE floating-point format, and their locations are encoded with $k\lceil\log_{2}(d)\rceil$ additional bits. We have $\omega=\frac{d}{k}-1$.
$\bullet$  Natural Compression (Horváth et al., 2022), a form of quantization in which floats are encoded into 9 bits instead of 32 bits. We have $\omega=\frac{1}{8}$.
$\bullet$  A combination of rand-$k$ and Natural Compression, in which the $k$ chosen values are encoded into 9 bits, which yields a total of $9k+k\lceil\log_{2}(d)\rceil$ bits. We have $\omega=\frac{9d}{8k}-1$.
$\bullet$  The $l_{1}$-selection compressor, defined as $C(x)=\text{sign}(x_{j})\|x\|_{1}e_{j}$, where $j$ is chosen randomly in $[d]$, with the probability of choosing $j^{\prime}\in[d]$ equal to $|x_{j^{\prime}}|/\|x\|_{1}$, and $e_{j}$ is the $j$-th standard unit basis vector in $\mathbb{R}^{d}$. $\text{sign}(x_{j})\|x\|_{1}$ is sent as a 32-bits float and the location of $j$ is indicated with $\lceil\log_{2}(d)\rceil$, so that this compressor communicates $32+\lceil\log_{2}(d)\rceil$ bits. Like with rand-$1$, we have $\omega=d-1$.
@@
We can see that LoCoDL, when combined with rand-$k$ and Natural Compression, converges faster than all other algorithms, with respect to the total number of communicated bits per client. We chose two different numbers $n$ of clients, one with $n<d$ and another one with $n>2d$, since the compressor of CompressedScaffnew is different in the two cases $n<2d$ and $n>2d$ (Condat et al., 2022a). LoCoDL outperforms CompressedScaffnew in both cases. As expected, all methods exhibit faster convergence with larger $n$. Remarkably, ADIANA, which has the best theoretical complexity for large $n$, improves upon DIANA but is not competitive with the LT-based methods CompressedScaffnew, 5GCS-CC, and LoCoDL. This illustrates the power of doubly-accelerated methods based on a successful combination of LT and CC. In this class, our new proposed LoCoDL algorithm shines. For all algorithms, we used the theoretical parameter values given in their available convergence results (Corollary 3.2 for LoCoDL). We tried to tune the parameter values, such as $k$ in rand-$k$ and the (average) number of local steps per round, but this only gave minor improvements. For instance, ADIANA in Figure 1 was a bit faster with the best value of $k=20$ than with $k=30$. Increasing the learning rate $\gamma$ led to inconsistent results, with sometimes divergence.
5 Conclusion
We have proposed LoCoDL, which combines a probabilistic Local Training mechanism similar to the one of Scaffnew and Communication Compression with a large class of unbiased compressors. This successful combination makes LoCoDL highly communication-efficient, with a doubly accelerated complexity with respect to the model dimension $d$ and the condition number of the functions. In practice, LoCoDL outperforms other algorithms, including ADIANA, which has an even better complexity in theory obtained from Nesterov acceleration and not Local Training. This again shows the relevance of the popular mechanism of Local Training, which has been widely adopted in Federated Learning. A venue for future work is to implement bidirectional compression (Liu et al., 2020; Philippenko & Dieuleveut, 2021; Dorfman et al., 2023). We will also investigate extensions of our method with calls to stochastic gradient estimates, with or without variance reduction, as well as partial participation. These two features have been proposed for Scaffnew in Malinovsky et al. (2022) and Condat et al. (2023), but they are challenging to combine with generic compression.
We have studied the strongly convex setting, because analyzing how the linear convergence rate depends on $d$, $\kappa$, $n$ provides valuable insights on the algorithmic mechanisms. It should be possible to derive sublinear convergence results in the general convex case, by studying the objective gap instead of the squared distance to the solution, as was done for CompressedScaffnew (Condat et al., 2022a) and TAMUNA (Condat et al., 2023). An analysis with nonconvex functions, however, would certainly require different proof techniques. In nonconvex settings, compression works well (Huang et al., 2022; Chen et al., 2024), but the properties of local training with variance reduction are less clear (Yi et al., 2024; Meinhardt et al., 2024).
@@ 5 Conclusion
Acknowledgments
This work was supported by funding from King Abdullah University of Science and Technology (KAUST): i) KAUST Baseline Research Scheme, ii) Center of Excellence for Generative AI, under award number 5940, iii) SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence.
@@ Acknowledgments
$=\left\|S\mathbf{\hat{x}}^{t}-\mathbf{x}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}$
$\quad+2n\mathbb{E}\!\left[\left\|\bar{d}^{t}-\mathbb{E}\!\left[\bar{d}^{t}\;|\;\mathcal{F}^{t},\theta=1\right]\right\|^{2}\;|\;\mathcal{F}^{t},\theta=1\right]$
@@
$\quad+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left\|\mathbf{u}^{t}-\mathbf{u}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}+\frac{p\chi}{\gamma}\left\|W\mathbf{\hat{x}}^{t}\right\|^{2}_{\boldsymbol{\mathcal{X}}}-2\langle\mathbf{u}^{t}-\mathbf{u}^{\star},\mathbf{\hat{x}}^{t}\rangle_{\boldsymbol{\mathcal{X}}}$
$=\frac{1}{\gamma}\left\|\mathbf{w}^{t}-\mathbf{w}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}+\frac{\gamma(1+2\omega)}{p^{2}\chi}\left(1-\frac{p^{2}\chi}{1+2\omega}\right)\left\|\mathbf{u}^{t}-\mathbf{u}^{\star}\right\|^{2}_{\boldsymbol{\mathcal{X}}}$
$\quad-\frac{p}{\gamma}\big{(}2\rho-\rho^{2}(1+\omega_{\mathrm{av}})-\chi\big{)}\left\|W\mathbf{\hat{x}}^{t}\right\|^{2}_{\boldsymbol{\mathcal{X}}}.$
@@
Appendix B Additional Experiments
The results with the ‘australian’ and ‘covtype.binary’ datasets from the LibSVM library, for the same logistic regression problem as in Section 4 with $\kappa=10^{4}$, are shown in Figures 4 and 5. Finally, we also run experiments on MNIST dataset (LeCun et al., 1998) in Figure 6. LoCoDL consistently outperforms the other algorithms in terms of communication efficiency.
Refer to caption Refer to caption
(a) $n=9$ (b) $n=41$
(c) $n=225$
Refer to caption Refer to caption
(a) $n=6$ (b) $n=100$
(c) $n=6,000$
Figure 5: comparison of LoCoDL and ADIANA using several compressors for logistic regression with the ‘covtype.binary’ dataset from the LibSVM, which has $d=54$ and 581,010 data points. We chose different values of $n$ to illustrate the three regimes $n<d$, $n>d$, $n>d^{2}$, as discussed at the end of Section 3.
Figure 6: Comparison of LoCoDL and ADIANA using several compressors for logistic regression. The task was classifying 7s and 8s from the MNIST (LeCun et al., 1998) dataset, which consists of $d=28\times 28=784$ dimensions and 14,118 data points. We chose $n=6$ (100$\times$ smaller than $d$).
Refer to caption Refer to caption
(a) $\alpha=1$ (b) $\alpha=10$
Figure 7: Comparison of LoCoDL and ADIANA with various compressors on logistic regression ($n=100$, $d=10$), with samples from the Dirichlet distribution of parameter $\alpha$, with $\alpha=1$ on the left and $\alpha=10$ on the right.
In an additional experiment, we investigate how heterogeneity of the functions influences the convergence. We consider logistic regression as above, but with synthetic data sampled from the Dirichlet distribution of parameter $\alpha$. If $\alpha$ is small, the Dirichlet distribution becomes similar to the uniform distribution over the simplex, which corresponds to the heterogeneous case where there is no similarity between the data. If $\alpha$ is large, the samples of the Dirichlet distribution tend to be similar to each other and concentrated around the middle point $(1/d,\ldots,1/d)$ of the simplex. We set $n=100$ and $d=10$, and a single random sample is assigned to each client. The results are shown in Figure 7, for $\alpha=1$ and $\alpha=10$. With these values of $n$ and $d$, ADIANA has a better theoretical complexity than LoCoDL. However, in practice, we observe that LoCoDL again outperforms ADIANA. For both methods, joint sparsification and quantization with rand-1 and natural compression performs best. There is no significant difference depending on the value of $\alpha$.